OpenStudy (anonymous):

Please explain to me the difference between gravitational potential and gravitational potential energy

OpenStudy (anonymous):

Gravitational potential is gravitational PE per unit mass. For example (on Earth): gravitational potential energy = PE = mgh gravitational potential = PE / m = gh

OpenStudy (anonymous):

i dont get the per unit mass thing, like really in simple words what is the difference or is there hardly anything

OpenStudy (anonymous):

At some height h, the gravitational potential is the amount of gravitational PE that a 1 kg mass would have. To find the potential, simply divide the PE by the mass. So say you have a 5 kg object 20 m above the ground. PE = mgh = (5kg)(20m)(g) = 981 J potential = PE / m = (981 J) / (5kg) = 196.2 J/kg It is important to realize, though, that potential exists EVERYWHERE. You do not need to have mass for there to be potential. On the other hand, potential energy does require the presence of mass (this is also shown by the formulas).